OpenStudy (anonymous):

X over 4 plus 8 = 16 x = -20 x = 4 x = 44 x = 11

OpenStudy (jackellyn):

x/4 + 8 = 16 x/4 = 16-8 x =4(16-8) Simplify for your answer.

OpenStudy (anonymous):

coo